# A Global Minimum Tax for Large Firms Only:

### v5

@@ 2.4 Tax regimes
Figure 2 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of $t_{M}$. The non-haven’s tax rates are given in panel (a), while the havens’ tax rates are in panel (b). The equilibrium regimes are summarized in panel (c). The solid curves give the tax rate(s) in the tax-competition equilibrium under the GMT; the dashed horizontal line is the unconstrained equilibrium tax rate $t_{i}^{0}$ (in Regime 0). Tax rates in both countries generally rise as the GMT increases, but there are distinct patterns in the different regimes. At the switch from Regime 1 to 2, the haven countries start splitting their tax rates and their tax rate on small MNEs drops discontinuously. At the switch from Regime 3 to 4, the non-haven splits its tax rate and discretely cut its tax rate for small MNEs to the level in the unconstrained Regime 0. The haven countries respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
3 Welfare effects of the GMT
Having fully described the tax-competition equilibrium in each regime, we can now turn to the welfare effects of the introduction of the GMT and of two gradual reforms: (i) a marginal increase in the GMT rate, $t_{M}$, and (ii) a marginal increase in the GMT coverage rate, $\phi$. As in Proposition 1, the qualitative nature of the propositions below holds independently of the marginal valuation of public funds, $\lambda$, and it includes revenue maximization ($\lambda\rightarrow\infty$) as a special case. Our discussion of the results will therefore focus on the implications that the GMT has on tax revenues.
3.1 Introduction of the GMT
We first compare welfare levels in the tax-competition equilibrium to those in the unconstrained equilibrium, Regime 0. This is summarized in:
Proposition 2
(GMT introduction)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1, and assume the GMT coverage rate is not too low so that $\phi\in(3/4,1)$. Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
  • (i)
    a rise in welfare in the non-haven country for all Regimes 1 to 4.
  • (ii)
    a rise in welfare in the haven country for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
  • (iii)
    a rise in world welfare for all Regimes 1 to 4.
Proof: See Appendix A.3.
Introducing a small GMT slightly higher than the havens’ unconstrained rate $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the haven countries. At the same time, profit shifting is reduced in comparison to the no-GMT benchmark due to a smaller tax differential. The sum of these effects must certainly benefit the non-haven. It also benefits the tax havens, as long as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits. However, once the GMT rate exceeds a certain threshold, $t_{M}^{+}$, which still lies in Regime 1, the reduced tax differential and thus the fall in shifted profits will dominate from the perspective of the haven countries. Therefore, for all levels of the GMT above $t_{M}^{+}$, tax havens lose from the introduction of the GMT while the non-haven gains. [^24] This conflict of interest does not arise in (Johannesen,, 2022, Proposition 6), where havens gain from the introduction of a GMT at any level, as long as profit shifting is not completely eliminated. In his setting where MNE shift profits only to the lowest-tax country, all the tax havens end up setting their tax rate to zero. Therefore, revenues in the unconstrained tax competition equilibrium are zero for each haven, and any positive GMT rate generates a revenue increase for the representative haven. See also footnote (19).
In the remaining Regimes 2 to 4, the tax differential for large MNEs is always smaller than the one in Regime 0. As long as large MNEs covered by the GMT account for a sufficiently high share of aggregate profits, $\phi\in(3/4,1)$, the havens lose from the introduction of $t_{M}$ higher than $t_{M}^{+}$ due to the reduced profit shifting of large MNEs. In this case, the effects of introducing the GMT on tax revenues and welfare in the non-haven and the haven countries remain opposed to each other throughout Regimes 2 to 4.252525If small MNEs not covered by the GMT account for a sufficiently large share of aggregate profits so that $\phi\in(0,3/4)$, then the introduction of the GMT rate will instead benefit the havens in Regime 3. See Appendix A.3 for details. Finally, the gains to the non-haven exceed the losses (if any) to the havens, so that the GMT introduction raises tax revenues and welfare worldwide in all regimes. These aggregate welfare gains result from the higher tax rates in both countries and the reduced levels of profit shifting.
3.2 Gradual reforms of the GMT
We now turn to two likely reform options after the GMT has been introduced. We first study the effects of a gradually increasing GMT rate. The welfare effects of this reform are summarized in:
Proposition 3
(Gradual reform: increasing GMT rate $t_{M}$)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1. A marginal increase in the GMT rate has the following effects:
  • (i)
    welfare in the non-haven country increases in all regimes, except at the regime-switching point $t_{M}^{1}$, where it discretely falls.
  • (ii)
    welfare in the haven country increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases for the remainder of Regime 1 and in Regime 2. It increases in Regime 3 and discretely falls at the last regime-switching point, $t_{M}^{3}$. It does not change in Regime 4.
  • (iii)
    welfare in the world increases in all regimes, except at the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, where it discretely falls.
Proof: See Appendices A.3 and A.4.
As is true for the introduction of the GMT (Proposition 2), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium. An exception is the regime-switching point $t_{M}^{1}$ where the havens start splitting their tax rate and set a discretely lower tax rate for small MNEs. Small MNEs then shift more profits to the havens, hurting the non-haven.
For the havens, the welfare (tax revenue) effects of an increase in the GMT depend on the negative effect of reduced profit-shifting opportunities and the positive effect of a higher tax rate in the non-haven. This leads to non-monotonic changes in the havens’ welfare. In Regime 1, the induced higher tax rate in the non-haven dominates initially, but as the GMT continues to rise, this is overcompensated by the reduced levels of profit shifting. The latter is also true throughout Regime 2. In Regime 3, the effect of a higher tax rate dominates again, as the non-haven commits to the uniform GMT rate in order to raise the havens’ tax rates. This regime, where commitment induces effective tax coordination, is similar to (the first part of) Regime 1. It ends at the regime-switching point $t_{M}^{3}$ where the non-haven starts splitting its tax rate. This leads to a discrete fall in the non-haven’s tax rate on small MNEs, which must hurt the tax havens.
The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive. At the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, however, where one of the countries starts splitting its tax rate and discretely lowers the tax rate on small MNEs, tax competition is tightened and world welfare falls at the margin.
Next we examine the welfare effects of another likely reform option: a marginal increase in the GMT coverage rate $\phi$, while keeping the GMT rate $t_{M}$ fixed. As shown in (8), this is achieved by reducing the threshold profit level $\pi_{M}$ above which the GMT rate must be charged. The results are summarized in:
Proposition 4
(Gradual reform: increasing GMT coverage rate $\phi$)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1. Except for special cases where a regime switch occurs, a marginal increase in the GMT coverage rate of $\phi\in(0,1)$ has the following effects:
  • (i)
  • (ii)
    welfare in the haven is unaffected in Regime 1 and decreases in Regimes 2 to 4.
  • (iii)
    world welfare is unaffected in Regime 1 and increases in Regimes 2 to 4.
Proof: See Appendix A.5.
In Regime 1, an increase in $\phi$ has no effect on welfare in any country, because all countries levy uniform tax rates. Changing $\phi$ does have effects, however, in the other regimes, where one or both countries split their tax rates. Since the equilibrium tax differential is always greater for small MNEs than for large MNEs, a higher $\phi$, which implies fewer small MNEs, reduces aggregate profit shifting. This unambiguously hurts the haven and benefits the non-haven. From the global perspective, a higher GMT coverage is always desirable, as it reduces pressure on tax competition for the profits of small MNEs.
There are two exceptional cases where increasing $\phi$ causes a regime switch. The first is at $t_{M}=t_{M}^{1}+\epsilon$ with $\epsilon$ being a small positive number, where a rise in $\phi$ pushes Regime 2 back to Regime 1 by weakening the havens’ incentive to split their tax rate. This regime switch benefits the non-haven and the world, while it has no effect on the havens. The other is at $t_{M}=t_{M}^{2}+\epsilon$, where a higher $\phi$ triggers a switch from Regime 3 to Regime 2, inducing the non-haven to set a tax rate above the GMT. This latter regime switch, however, leaves welfare in all countries and the world unaffected.
@@ 4 Quantitative implications
We now examine which regime can be expected under the currently imposed GMT rate. For this purpose, we use a calibrated version of our model and explore its quantitative effects. We calibrate our unconstrained model without GMT (Regime 0) to match basic data on international profit shifting. The results are summarized in Table 2. The coverage rate of the GMT is set at $\phi=0.9$, in line with OECD, 2020a and our own calculations from the ORBIS database (see Figure 1 and Appendix B). The marginal valuation of public funds (MVPF) is set at $\lambda=1.5$ in our calibration benchmark, corresponding to the benchmark value used in Hebous and Keen, (2023). The figures for total pre-tax profits of MNEs, $\Pi$, and the number of tax havens, $H$, are taken from Tørsløv et al., (2023) and the Country-by-Country Report Statistics (CbCR); see the notes to Table 2 for details. We then calculate the cost parameter of profit shifting, $\delta$, to exactly match the GDP-weighted average of effective tax rates in non-haven countries of $18.6\%$, as reported in Tørsløv et al., (2023). Among the non-targeted moments, our model somewhat underestimates the havens’ tax rate in the data, as well as the total profits shifted by MNEs and –accordingly– the non-havens’ loss of tax revenues. Overall, however, our simple model explains the data reasonably well.
@@
$\Pi=4,623$bUSD Total pre-tax profits of MNEs Tørsløv et al., (2023), CbCR
$\phi=0.9$ Profit share of MNEs covered by GMT OECD, 2020a , Orbis
@@
$\lambda=1.5$ Marginal valuation of public funds Hebous and Keen, (2023)
$\delta=27.4$ Cost of profit shifting Calibrated
@@
Corp. tax rate of non-haven (Regime 0): $t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}$ $18.6\%$ $18.6\%$ Tørsløv et al., (2023)
@@
Corp. tax rate of haven (Regime 0): $t_{h}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ $9.3\%$ $13.7\%$ Tørsløv et al., (2023)
MNE’s shifted profits: $\sum_{h}\theta_{h}^{0}\Pi$ $660$bUSD $969$bUSD Tørsløv et al., (2023)
Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ $129$bUSD $247$bUSD Tørsløv et al., (2023)
Table 2: Calibration of the model
@@
Notes: Data are from 2019. The definitions of non-haven and haven countries follow Tørsløv et al., (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ”Table1” and ”Table U1” in the excel file ”1975-2019 updated estimates: Tables”). The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world. There are $40$ haven countries across the world. The corporate tax rates for the representative non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates. Tørsløv et al., (2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7). Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs. With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
Figure 3: Equilibrium welfare and shifted profits
Notes: The figure shows the equilibrium levels of welfare normalized by the MVPF, $\lambda$, in the non-haven (a) and the sum of the $H=40$ haven countries (b), and the aggregate shifted profits (c), for different rates of GMT. Parameter values are given in Table 2. The solid curves in each panel are the equilibrium values under the partial GMT ($\phi=0.9$); the dashed horizontal lines are those in the unconstrained equilibrium (Regime 0).
For this calibrated version of our model, the effects of introducing a GMT at varying tax rates $t_{M}$ on equilibrium tax schedules are shown in Figures 2, which we already discussed in Section 2.4. Figure 3 illustrates the effects on welfare in both non-haven and haven countries, and those on aggregate shifted profits, as compared to their counterparts in the unconstrained equilibrium (Regime 0). To make numbers comparable with tax revenues, the levels of welfare are normalized by the MVPF, $\lambda$.
The most important result is that the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 2. In this regime, the havens find it profitable to split their tax rate and to set a lower tax rate for small MNEs as compared to large MNEs (Proposition 1(ii)). [^26] This result also holds if we assume that the first stage decision of whether to split tax rates or not is made by each tax haven individually (see Appendix A.2). In this case the range of GMT rates that support Regime 1 will be smaller than in our benchmark model, while the range of $t_{M}$ that supports an equilibrium in Regime 2 rises accordingly. This corresponds to the observations mentioned in the Introduction that several tax havens have announced to maintain their low statutory tax rate while topping up the taxation of large MNEs to comply with the GMT, known as the QDMTT (see footnote 5). The GMT introduction reduces the total amount of profits shifted, thereby increasing the non-haven’s welfare while reducing the havens’ welfare (tax revenues).
Panel (a) of Table 3 further quantifies the welfare and revenue effects of a 15% GMT rate for our calibrated model, with the benchmark MVPF of $\lambda=1.5$. Welfare and tax revenue increases in the non-haven country clearly outweigh the revenue losses in the havens, leading to an increase in global welfare and tax revenues. Our calibrated gain in tax revenue in the non-haven (107 billion USD) is close to the summed revenue gains for Europe and the U.S. in Baraké et al., (2022) and worldwide tax revenue gains are in the same range as in Hugger et al., (2024) (see footnote 2). [^27] The percentage increases in corporate tax revenue in our model are not comparable to those in Baraké et al., (2022) and Hugger et al., (2024), as we do not model purely domestic firms. The welfare gain for the non-haven is smaller than its revenue gain, because the private income in the non-haven (the post-tax profits of MNEs net of profit-shifting costs) falls after the introduction of GMT.
(a) MVPF: $\lambda=1.5$.  Regime 0 to 2 Non-haven Haven World
Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
@@
Tax rate after GMT introduction $19.9\%$ for all $\begin{cases}9.9\%&\text{for small}\\ 15.0\%&\text{for large}\end{cases}$ —–
Welfare change in bUSD (change) $53$ ($1.6\%$) $-7$ ($-12.1\%$) $46$ ($1.4\%$)
Revenue change in bUSD (change) $107$ ($14.5\%$) $-7$ ($-12.1\%$) $99$ ($12.4\%$)
Shifted-profits change in bUSD (change) $267$ $(41.9\%)$ $-267$ $(-41.9\%)$ 00 $(0\%)$
(b) MVPF: $\lambda=3$.  Regime 0 to 1 Non-haven Haven World
Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
Tax rate after GMT introduction $20.9\%$ for all $15.0\%$ for all —–
Welfare change in bUSD (change) $122$ ($6.4\%$) $2$ ($2.0\%$) $124$ ($6.1\%$)
Revenue change in bUSD (change) $168$ ($26.0\%$) $2$ ($2.0\%$) $170$ ($22.6\%$)
Shifting-profits change in bUSD (change) $425$ $(36.8\%)$ $-425$ $(-36.8\%)$ 00 $(0\%)$
(c) MVPF: $\lambda=10$.  Regime 0 to 1 Non-haven Haven World
Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
Tax rate after GMT introduction $21.3\%$ for all $15.0\%$ for all —–
Welfare change in bUSD (change) $169$ ($17.3\%$) $12.4$ ($9.3\%$) $181$ ($16.3\%$)
Revenue change in bUSD (change) $185$ ($31.1\%$) $12.4$ ($9.3\%$) $197$ ($27.1\%$)
Shifting-profits change in bUSD (change) $463$ $(32.2\%)$ $-463$ $(-32.2\%)$ 00 $(0\%)$
Table 3: Quantitative effects of introducing a $15\%$ GMT
Notes: Each of panels (a) to (c) show the effects of a $15\%$ GMT introduction for different values of the MVPF, $\lambda\in\{1.5,3,10\}$. Results are based on the calibrated parameter values described in Table 2; the GMT coverage rate is $90\%$ ($\phi=0.9$) in particular. The welfare levels are normalized by the MVPF, $\lambda$. With $\lambda=1.5$, the $15\%$ GMT introduction leads the economy from Regime 0 to 2; with $\lambda\in\{3,10\}$, it leads instead to Regime 1. The haven refers to the sum of the $H=40$ haven-countries. Due to rounding numbers, the sum of the welfare/revenue change of the non-haven and the haven in billion USD is not necessarily equal to that of the world.
One may argue that the MVPF is indeed higher than our benchmark value of $\lambda=1.5$, as evidenced by the very substantial efforts, discussed in the introduction, that non-haven countries take to increase their corporate tax revenues. Panels (b) and (c) of Table 3 therefore present the results of two robustness checks with high values of the MVPF: $\lambda=3$ and $\lambda=10$. The latter value, given in Panel (c) approaches a Leviathan government objective of tax revenue maximization. With these increased values of $\lambda$, the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 1, in which both countries employ a uniform tax rate on all MNEs. With the higher valuation of tax revenues in the non-haven country, the tax havens will find it profitable to commit to a uniform GMT rate for all MNEs, which in turn leads the non-haven to choose a higher uniform tax rate above the GMT, as compared to the benchmark in Panel (a). The GMT rate of $15\%$ therefore acts as a coordination device between the non-haven and the havens, benefitting both countries. [^28] These results depend on the assumption that tax havens make their first stage choice collectively. With decentralized first stage decisions of tax havens (cf. Appendix A.2), the upper boundary for a Regime 1 equilibrium, $t_{M}=t_{M}^{a}$, is calculated to be $9.37\%$, $9.42\%$, and $9.45\%$ for respective values of the marginal valuation of public funds, $\lambda=1.5$, $3$, and $10$. This implies that the introduction of a $15\%$ GMT rate leads to a Regime 2 equilibrium regardless of the level of $\lambda$. For $\lambda=1.5$, the quantitative effects under decentralized decision making of havens are the same as those in Panel (a) of Table 3. For $\lambda=3$ and $\lambda=10$, the effects differ from those in Panels (b) and (c) of Table 3, however, and yield a conflict of interests between the non-haven and the havens that is analogous to the one in Panel (a). Although observations at the time of writing somewhat favor the GMT resulting in split tax rates in the havens, corresponding to Regime 2, we leave open the question of which regime will eventually emerge.
(a) Gains from $t_{M}=0.15$ to $t_{M}=0.18$ Non-haven Haven World
Welfare change in bUSD (change) $34$ ($1.0\%$) $-17$ ($-31.0\%$) $18$ ($0.5\%$)
Revenue change in bUSD (change) $59$ ($7.0\%$) $-17$ ($-31.0\%$) $42$ ($4.7\%$)
Shifting-profits change in bUSD (change) $146$ $(37.9\%)$ $-146$ $(-37.9\%)$ 00 $(0\%)$
(b) Gains from $\phi=0.9$ to $\phi=1.0$ Non-haven Haven World
Welfare change in bUSD (change) $5.4$ ($0.2\%$) $-0.5$ ($-0.9\%$) $4.9$ ($0.1\%$)
Revenue change in bUSD (change) $10.7$ ($1.3\%$) $-0.5$ ($-0.9\%$) $10.3$ ($1.1\%$)
Shifting-profits change in bUSD (change) $27$ $(7.0\%)$ $-27$ $(-7.0\%)$ 00 $(0\%)$
Table 4: Quantitative effects of gradual reforms
@@
Notes: Panel (a) shows gains/losses from an increase in the GMT rate from $t_{M}=0.15$ to $0.2$; panel (b) shows gains/losses from an increase in the GMT coverage rate from $\phi=0.9$ to $1.0$. Results are based on the calibrated parameter values described in Table 2; the MVPF is $\lambda=1.5$ in particular. The welfare levels are normalized by the MVPF, $\lambda$. The haven refers to the sum of the $H=40$ haven-countries. Due to rounding numbers, the sum of the welfare/revenue change of the non-haven and the haven in billion USD is not necessarily equal to that of the world.
Finally, Table 4 provides a quantitative assessment of two gradual reforms that are likely to follow the newly introduced GMT rate of $15\%$ with incomplete coverage. The first reform is to increase the GMT rate from $15\%$ to $18\%$ while maintaining the coverage rate at $\phi=90\%$ (Panel (a)). The second is to broaden the GMT coverage from $90\%$ to $100\%$ while keeping the GMT rate at $15\%$ (Panel (b)). In both scenarios, we assume our benchmark MVPF of $\lambda=1.5$ and thus, the equilibrium remains in Regime 2 after the reforms. The results indicate that the non-haven gains more from the $18\%$ GMT rate ($+1.0\%$ in welfare; $+7.0\%$ in revenues) than from the full GMT coverage ($+0.2\%$ in welfare; $+1.3\%$ in revenues). The greater gains for the non-haven from the $18\%$ GMT rate are driven by the stronger reduction in profit shifting to the havens ($-37.9\%$), as compared to a full GMT coverage ($-7.0\%$). Hence the increase in the GMT rate creates a much sharper conflict of interest with the tax havens, as compared to the extension of GMT coverage. In this sense, covering more MNEs under the GMT may be a politically more feasible option than increasing the GMT rate. However, even a full coverage of the GMT will likely bring only moderate gains to the non-haven and to the world, while entailing higher administrative and compliance costs.

### v6

@@ 2.1 Tax regimes
Figure 1 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of the GMT rate $t_{M}$. The non-haven’s tax rates are given in panel (a), while the havens’ tax rates are in panel (b). The equilibrium regimes are summarised in panel (c). The solid (dotted) curves are the tax rates for large MNEs (small MNEs) in the tax-competition equilibrium under the GMT; the dashed horizontal lines are the unconstrained equilibrium tax rate $t_{i}^{0}$ (in Regime 0). Tax rates in both countries generally rise as $t_{M}$ increases, but there are distinct patterns in the different regimes. At the switch from Regime 1 to 2, the havens begin to split their tax rates, and the rate for small MNEs falls discontinuously. At the switch from Regime 3 to 4, the non-haven splits its tax rate and discretely cut its tax rate for small MNEs to the level in the unconstrained Regime 0. The havens respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
2.2 Welfare effects of the GMT
We briefly summarise how the GMT introduction affects national and global welfare. Introducing a GMT rate $t_{M}$ slightly above the havens’ unconstrained one $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the havens. At the same time, profit shifting is reduced in comparison to the unconstrained Regime 0 due to a smaller tax differential. The sum of these effects must benefit the non-haven. It also benefits the havens, as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits.
Once $t_{M}$ exceeds a certain threshold within Regime 1, the relationship reverses for the havens, as the loss of shifted profits dominates. Introducing any GMT rate above this threshold benefits the non-haven but harms the havens, with the net effect raising global welfare. Supplemental Appendix C provides a formal proposition and proof of these results.
We also study the effects of marginal increases in the GMT tax rate $t_{M}$ and the GMT coverage rate $\phi$. An important result is that welfare in the non-haven and the world increases with $t_{M}$ in general, and the welfare gains become even larger as $\phi$ increases. A higher $\phi$ reduces the effectiveness of the split tax system and limits tax competition for small MNEs, thereby amplifying the welfare gains from raising $t_{M}$. See Supplemental Appendices D and E for formal discussions and Appendix F for graphical illustrations.
@@ 3 Quantitative implications
We now explore the quantitative importance of our theoretical results. We modify the model of the previous section by assuming that the havens decide tax policies in a decentralised manner (see footnote 16 and Supplemental Appendix B). The decentralised decision making leads to competition between the havens and thus makes it more likely that they split their tax rates. [^19] The new regime switching rate dividing Regimes 1 and 2 is smaller ($t_{M}^{b}<t_{M}^{1}$), and thus Regime 1 is less likely to emerge, as $t_{M}^{0},t_{M}^{b})\subset[t_{M}^{0},t_{M}^{1})$. This is because the commitment to a single GMT rate by a single haven does not induce the non-haven to raise its tax rate as much as the collective commitment by multiple havens, making gains from commitment smaller.
We calibrate the unconstrained model with no GMT (Regime 0) to match international profit-shifting data, as summarised in Table 2. The GMT coverage rate is set to $\phi=0.9$, following OECD, (2020) and our calculations from Orbis. [^20] See Supplemental Appendix A for details. The number of tax havens is set to $H=40$, as in Tørsløv et al., (2023). Combining their data with Country-by-Country Reporting (CbCR) yields total MNE pre-tax profits of $\Pi=4,623$ billion USD. The marginal value of public funds (MVPF), $\lambda$, and the cost parameter of profit shifting, $\delta$, are jointly calibrated to minimise the squared distance between the model-predicted non-haven tax rate and shifted profit share and their empirical counterparts. For non-targeted moments, our model slightly underestimates the havens’ tax rate, $t_{h}^{0}$, and the non-haven’s revenue losses, $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$, but overall fits the data reasonably well.
Table 2. Calibration of the Model.
@@
$\phi=0.9$ Profit share of MNEs covered by GMT OECD, (2020), Orbis
@@
$\Pi=4,623$bUSD Pre-tax profits of MNEs Tørsløv et al., (2023), CbCR
$\lambda=2.12$ Marginal value of public funds Calibrated
$\delta=17.80$ Cost of profit shifting Calibrated
@@
Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al., (2023)
Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al., (2023), CbCR
@@
Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al., (2023)
Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al., (2023)
@@
Notes: Data are from 2019. The definitions of non-haven and haven countries follow Tørsløv et al., (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables’). The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world. There are $40$ haven countries across the world. The corporate tax rates for the representative non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates. Tørsløv et al., (2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7). Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs. With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
Using this calibrated version of our model, we examine the quantitative effects of GMT reforms, as summarised in Table 3. The key result is that a 15% GMT rate with partial coverage ($\phi=0.9$) yields the equilibrium in Regime 2 (Table 3(a)): the havens split their tax rates and set a lower tax rate ($10.1\%$) for small MNEs as compared to large MNEs ($15.0\%$) (Proposition 1(ii)). [^21] Above the regime switching rate of $t_{M}^{b}=0.094$, Regime 2 prevails. This result aligns with observations that low-tax countries like Ireland plan to retain their base rate (12.5% in Ireland’s case) while applying top-up taxes on large MNEs under the GMT (QDMTT).
Table 3. Quantitative Effects of GMT Reforms.
(a) Gains from a 15% GMT under $\phi=0.9$ Non-haven Haven World
@@
Tax rate after GMT introduction $20.3\%$ for all $\begin{cases}10.1\%&\text{for small}\\ 15.0\%&\text{for large}\end{cases}$ —–
Welfare change in bUSD (% change) $87.2$ ($3.5\%$) 3.8-3.8 ($4.3\%$) $83.4$ ($3.2\%$)
Revenue change in bUSD (% change) $139.5$ ($20.6\%$) 3.8-3.8 ($4.3\%$) $135.7$ ($17.7\%$)
Profit change in bUSD (% change) $359.9$ $(37.5\%)$ 359.9-359.9 $(-37.5\%)$ 0 $(0\%)$
(b) Gains from a 15% GMT under $\phi=1.0$ Non-haven Haven World
Tax rate after GMT introduction $20.6\%$ for all $15\%$ for all —–
Welfare change in bUSD (% change) $95.6$ ($3.8\%$) 3.6-3.6 ($4.1\%$) $92.0$ ($3.6\%$)
Revenue change in bUSD (% change) $153.0$ ($22.6\%$) 3.6-3.6 ($4.1\%$) $149.3$ ($19.5\%$)
Profit change in bUSD (% change) $393.0$ $(41.0\%)$ 393.0-393.0 $(-41.0\%)$ 0 $(0\%)$
@@
Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$) and panel (b), from $t_{M}=0$ to $0.15$ with full coverage $\phi=1.0$. The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven. Results are based on the calibrated parameter values described in Table 2. The welfare levels are normalised by the MVPF, $\lambda$. The ‘haven’ here refers to the sum of the $H=40$ haven countries.
The non-haven’s revenue rises due to weaker tax competition (higher tax rates) and reduced profit shifting, while the havens’ revenues fall as the loss of shifted profits outweighs the gains from milder tax competition. The calibrated revenue gain in the non-haven (139.5 bUSD) is close to the combined gains for Europe and the US in Baraké et al., (2022, Table 1) ($67+58=125$ bEUR), and the global revenue gain (135.7 bUSD) is comparable to Hugger et al., (2024, Section 8.2) (155–192 bUSD). Since falling shifted profits reduce MNEs’ private benefits, the non-haven’s welfare rises less than their revenues. [^22] In the havens without MNE headquarters, revenues coincide with welfare.
Do the quantitative effects of introducing GMT change if the coverage is extended? The effects of $t_{M}=0.15$ with full coverage ($\phi=1.0$) yields the non-splitting equilibrium in Regime 1 (Table 3(b)). By eliminating competition for small MNEs, the extended coverage increases revenue and welfare gains worldwide by 13.6 bUSD and 8.6 bUSD, respectively. Based on the estimates by Gaul et al., (2022), these costs would amount to about 10.5 bUSD for our sample, implying a small net revenue gain (+3.1+3.1 bUSD) but a net welfare loss (1.9-1.9 bUSD) globally. [^23] Using a survey of German MNEs (see Becker et al., , 2012 for details on their characteristics), Gaul et al., (2022) estimate the average one-time expenses of implementing GMT per firm at 1.6 million EUR. Applying this to the number of small MNEs not covered by GMT in our database ($\approx$6,000; see Table A.1 in Supplemental Appendix A), the total compliance costs of newly covered MNEs are 1.6 mEUR $\times$ 6,000 $\times$ 1.1 USD/EUR $\approx$ 10.5 bUSD. Clifford et al., (2025) also estimate size-dependent compliance costs for German MNEs and similarly find that an increased coverage of the GMT would provide only limited gains for Germany.